THE RATIO OF THE INTENSITIES AT MINIMA TO MAXIMA IN THE YOUNGS DOUBLE SLIT EXPERIMENT IS 9 :25.FIND THE RATIO OF THE WIDTHS OF THE TWO SLITS


1
Expert's answer
2020-08-02T15:03:56-0400
"\frac{I_{min}}{I_{max}}=\frac{(a_1-a_2)^2}{(a_1+a_2)^2}=\frac{9}{25}\\\frac{a_1-a_2}{a_1+a_2}=\frac{3}{5}\\\frac{a_1}{a_2}=\frac{4}{1}\\\frac{w_1}{w_2}=\frac{a_1^2}{a_2^2}=\frac{16}{1}"